# Surf-D: High-Quality Surface Generation for Arbitrary Topologies using Diffusion Models

### 28 Nov 2023 v1

@@ 4.5 Latent Interpolation
Our point-based auto-encoder provides a compact latent space, which enables the latent diffusion framework to learn the semantics with high efficiency. In Fig. 9, we show different shapes generated by the naive linear interpolation in the latent space. Even without additional semantic annotation during training, the shape can change smoothly during the latent code interpolation, which demonstrates continuous and high-fidelity shape embedding.
@@
Figure 9: Qualitative results of latent code interpolation. We sample latent codes $z_{1}$ and $z_{2}$ from latent space and decode the linear interpolated latent codes.
@@ 5 Conclusion
The appendix covers: more experimental results (Sec. A); ablation study (Sec. B); implementation details (Sec.C); and limitations and future work. (Sec. D).
Appendix A More Results
Finally, we show that the generated clothing can be easily integrated into a simulation platform***https://www.marvelousdesigner.com/ for virtual try-on, which is shown in Fig. 12. These compelling results validate our high-quality 3D asset generation.
Appendix B Ablation Study
B.1 Ablation of Curriculum Learning Strategy
In Fig. 14, we present the qualitative results of an ablation study on curriculum learning strategy, where the learning strategy promotes a higher-quality geometry.
[…]
Figure 12: Garment Generation for Virtual Try-on . The generated garments can seamlessly integrate into an animation pipeline for virtual try-on.
@@ B.1 Ablation of Curriculum Learning Strategy
Figure 13: Comparison of Loss Curve.
@@
B.2 Ablation of the Training Loss Function
@@
[…]
Figure 15: Comparison of different reconstruction loss.
Appendix C Implementation Details
@@
C.1 Query points sampling strategy for UDF decoder
@@ C.2 Hyperparameters settings
Appendix D Limitations and Future work
Although Surf-D demonstrates superior performance in shape generation across multiple modalities, it has some limitations. As the failed case shown in Fig. 16, Surf-D is unable to accurately reconstruct the iron chain (see the zoom-in area). This is because the frequency of geometric details in the iron chain portion is higher than the resolution of the grid for the marching cube, resulting in the loss of geometric information in that part. It would be valuable to explore how to extract mesh from a 3D UDF field efficiently and faithfully. As current mesh extraction methods from the UDF field still require gradients, extracting mesh from the UDF field in a gradient-free way will be an interesting direction. Furthermore, our model requires a relatively large amount of expensive 3D data for capturing the distribution of versatile shapes. Thus, learning shape knowledge from 2D images would significantly alleviate the need for 3D data, which could also be an intriguing research direction.
[…]
Figure 16: Failure case of Surf-D. Surf-D fails to represent the tiny iron chain.

### 23 Mar 2024 v2

@@ 4.6 Latent Interpolation
Figure 11: Qualitative results of latent code interpolation.
Our point-based auto-encoder provides a compact latent space, which enables the latent diffusion framework to learn the semantics with high efficiency. In Fig. 11, we sample latent codes $z_{1}$ and $z_{2}$ from latent space and decode the linear interpolated latent codes to show different shapes generated by the naive linear interpolation in the latent space. Even without additional semantic annotation during training, the shape can change smoothly during the latent code interpolation, which demonstrates continuous and high-fidelity shape embedding.
@@ 5 Conclusion
[…]
Figure 12: Failure case of Surf-D.
Limitations and Future Works Although Surf-D demonstrates superior performance in shape generation across multiple modalities, it has some limitations. As shown in Fig. 12, Surf-D is unable to accurately reconstruct the iron chain (see the zoom-in area). This is because the frequency of geometric details in the iron chain portion is higher than the resolution of the grid for the marching cube, resulting in the loss of geometric details in that part. It would be valuable to explore how to extract mesh from a UDF field efficiently and faithfully. As current mesh extraction methods from the UDF field still require gradients, extracting mesh from the UDF field in a gradient-free way will be an interesting direction. Furthermore, our model requires a relatively large amount of expensive 3D data for capturing the distribution of versatile shapes. Thus, learning shape knowledge from 2D images would significantly alleviate the need for 3D data, which could also be an intriguing research direction.
@@
This supplementary material covers: more experimental results (Sec. 0.A); additional ablation studies (Sec. 0.B); implementation details (Sec.0.C).
[…]
Figure 13: Comparison of learning SDFs. Our point-based AutoEncoder learns continuous SDFs and produces higher-quality surfaces.
@@
Appendix 0.A More Experimental Results
0.A.1 Grid-based Distance Fields V.S. Continuous Distance Fields on SDFs
In Sec. 4.5 of the main paper, we demonstrated the effectiveness of our point-based AutoEncoder design for shape embedding of UDFs compared to the grid-based approach [12]. Here, we further highlight the advantages of continuous distance fields in learning SDFs.
Specifically, we follow SDFusion [12] to first make the car objects in ShapeNet [6] dataset watertight, and compare our point-based AutoEncoder with grid-based AutoEncoder used in SDFusion for learning SDFs. As shown in Fig. 13, our pipeline can be adjusted to learn continuous signed distance field and produce higher-quality surfaces than the previous method.
0.A.2 More Results on Text-to-Shape Generation
@@ 0.A.3 More Results of Unconditional Shape Generation
Appendix 0.B Additional Ablation Study
0.B.1 Ablation of Curriculum Learning Strategy
@@
Figure 16: Loss Curve Comparison.
Figure 17: Ablation results of curriculum learning strategy.
[…]
@@
Figure 18: Comparison of different reconstruction loss.
In Fig. 18, we present the qualitative results of an ablation study on curriculum learning strategy under the same training setting, where the learning strategy promotes a higher-quality geometry.
0.B.2 Ablation of the Training Loss Function
@@
Appendix 0.C Implementation Details
@@
Query points sampling strategy for UDF decoder
@@
Dataset size and training time
Deepfashion3D [79] contains $563$ diverse cloth items of $9$ categories. Pix3D [62] contains $395$ shapes of $9$ categories with $10069$ corresponding images. Text-to-shape contains $75,344$ text descriptions for around $10$k chair and table objects. For the training, we use a single 40G A100 GPU. The training time differs on datasets. For Deepfashion3D [79] and Pix3D [62] dataset, it takes $3$ days to train the AutoEncoder and $2$ days to train the diffusion model. For Text-to-shape in ShapeNet [6] dataset, it takes $5$ days to train the AutoEncoder and $3$ days to train the diffusion model.